OpenStudy (anonymous):

hey can pls anyone solve this question integration sign x/x^2+1 dx

OpenStudy (anonymous):

now, \[x \div(x ^{2}+1)=2x \div2(x ^{2}+1)\]

OpenStudy (anonymous):

ln(t+1)/2+c , wer t=x^2 :)

OpenStudy (anonymous):

got it ?

OpenStudy (anonymous):

let \[x ^{2}+1 = z\] or, \[2xdx = dz\] now, \[\int\limits_{}^{} [x \div (x ^{2} + 1)] dx\] = \[\int\limits_{}^{} dz/z\] = \[\ln z + c\] = \[\ln (x ^{2}+1) + c\]

OpenStudy (anonymous):

@ toriqur : well done , but u missed a 2 while substituting :) xdx=dz/2 !

OpenStudy (anonymous):

oh, sorry. really i am. just a mistake, bro.

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

i still didnt get it :(